OpenStudy (bimbim2697):

a rectangle has a length of 6x^3 y^2 - 1 and a width of 3xy+2.The formaula for the perimeter of a rectangle is P=2l+2w ,where l is length and w is the width.What is the perimeter of the rectangle?Simplify your answer.

hero (hero):

P = 2l + 2w P = 2(l + w) P = 2(6x^3y^2 - 1 + 3xy + 2) P = 12x^3y^2 - 2 + 6xy + 4 P = 12x^3y^2 + 6xy + 2 P = 2(x^3y^2 + 3xy + 1)

OpenStudy (anonymous):

P=2 ( l+ w ) = 2 [ (6x^3 y^2 - 1 ) + (3xy+2)] = 2 ( 6x^3 y^2 + 3xy + 1)

OpenStudy (anonymous):

There's nothing to simplify or calculate !

hero (hero):

I hate when I still make a mistake regardless of posting solution first.

hero (hero):

P = 2l + 2w P = 2(l + w) P = 2(6x^3y^2 - 1 + 3xy + 2) P = 12x^3y^2 - 2 + 6xy + 4 P = 12x^3y^2 + 6xy + 2 P = 2(6x^3y^2 + 3xy + 1)